OpenStudy (unicwaan):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = the quantity x minus seven divided by the quantity x plus three. and g(x) = quantity negative three x minus seven divided by quantity x minus one. I understood how to do the first one but got lost on the second one, can someone help me?

OpenStudy (unicwaan):

\[f(x)=\frac{ x-7 }{ x+3 }\] \[g(x)=\frac{ -3x-7 }{ x-1 }\]